OpenStudy (anonymous):

solve the equation or formula for the indicated variable s=5r^2t, for t

OpenStudy (anonymous):

solve the equation or formula for the indicated variable s=5r^2t, for t

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

ln ( x^r ) = r * ln(x) since rule of logs lets us bring out powers, so, lets take the ln of each side, you should get... ln( S ) = ln ( 5r^4t) remember rules of logs lets us bring out powers, sooo... ln( S ) = 4t * ln( 5r ) now divide both sides by the ln( 5r ) [ ln( S ) / ln( 5r ) ] = 4t now lets divide both sides by 4 to get t by itself, we get... [ ln( S ) / ln( 5r ) ] / 4 = t Hopefully this helped

OpenStudy (anonymous):

i need a medal

OpenStudy (anonymous):

thanxs

OpenStudy (anonymous):

always